OpenStudy (anonymous):

Medal and fan Given the following functions f(x) and g(x), solve f[g(10)] and select the correct answer below: f(x) = 10x + 8 g(x) = x + 9 2,052 98 190 198

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@welshfella

OpenStudy (igreen):

First, solve g(10)..which is just plugging in 10 for \(\sf x\).

OpenStudy (igreen):

\(\sf g(x) = x + 9\) \(\sf g(10) = 10 + 9\)

OpenStudy (igreen):

What do you get?

OpenStudy (anonymous):

-9 and 0??

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

-9?

OpenStudy (igreen):

\(\sf 10 + 9 \ne -9~or~0\)

OpenStudy (anonymous):

idk, i'm confused on this

OpenStudy (igreen):

You can't add \(\sf 10 + 9\)?

OpenStudy (anonymous):

it's 19, i thought there was more to it

OpenStudy (igreen):

g(10) means we plug in 10 for 'x' in g(x) = x + 9.

OpenStudy (igreen):

Which gives us: \(\sf g(10) = 10 + 9\) Simplify: \(\sf g(10) = 19\) You're correct..now we have: \(\sf f(19)\) Plug in \(\sf 19\) for 'x' in \(\sf f(x) = 10x + 8\) now.

OpenStudy (igreen):

Understand?

OpenStudy (anonymous):

f(19) = 10(19) + 8? Or is it f(x) = 10x + 8

OpenStudy (anonymous):

f(19) = 10x + 8*

OpenStudy (anonymous):

Help

OpenStudy (igreen):

f(19) = 10(19) + 8 is correct..

OpenStudy (igreen):

\(\sf f(x) = 10x + 8\) \(\sf f(19) = 10(19) + 8\)

OpenStudy (igreen):

Now simplify the right side

OpenStudy (anonymous):

198 is what i got

OpenStudy (igreen):

That is correct.

OpenStudy (anonymous):

Thanks, could you help me with a few more :D?

OpenStudy (igreen):

Sure, just close this and open a new one.

OpenStudy (anonymous):

Alright